### v3

# Is UEFA really committed to sporting merit? A misaligned rule in the post-2024 Champions League

Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification of the post-2024 UEFA Champions League is not always based on sporting merit. Consequently, a match with misaligned incentives has been barely avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended to safeguard the integrity of the competition. Fairness can be easily guaranteed by reversing the order of two articles in the Champions League regulation. Since UEFA repeatedly commits the same mistake that is widely known among researchers, governing bodies in sports are called to work more closely together with the scientific community.
Keywords: fairness; incentive compatibility; integrity; tournament design; UEFA Champions League
MSC class: 62F07, 90B90, 91B14
JEL classification number: C44, D71, Z20
Hoc est, quod ait Heraclitus: “In idem flumen bis descendimus et non descendimus.111 Source: https://www.perseus.tufts.edu/hopper/text?doc=Perseus%3Atext%3A2007.01.0080%3Aletter%3D58
[…]
1 Introduction
Ensuring the fairness and integrity of competitions is a key responsibility of sports policy (Sugden et al.,, 2023). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League with the following words (UEFA, 2024a, ):
UEFA has clearly shown that we are fully committed to respecting the fundamental values of sport and to defending the key principle of open competitions, with qualification based on sporting merit, fully in line with the values and solidarity-based European sports model.
However, is the UEFA Champions League qualification based on sporting merit under any circumstances? Is there any issue in its rules that compromises the fairness of this tournament and can potentially jeopardise competition?
The current note uncovers that, while UEFA may be committed to sporting merit in principles, their rules do not avoid an unfair game in a domestic league. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots are assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig.
This problem is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, its solution has been known since March 2013 when the first version of Dagaev and Sonin, (2018) become available.
2 Related literature
Since tanking—the act of deliberately losing a game—is against the spirit of sports, incentive compatibility is an extensively discussed topic in the field of tournament design (Kendall and Lenten,, 2017; Csató,, 2021). Probably the most serious type of incentive incompatibility is when a team can be strictly better off by losing. According to our knowledge, such a possibility has been identified first by Dagaev and Sonin, (2018) in the UEFA Europa League. Several qualification tournaments for the FIFA World Cups and the UEFA European Championships have been verified to suffer from the same weakness (Csató, 2020a, ), even allowing for a match where the dominant strategy of both opponents is to play a draw (Csató, 2020c, ).
In the past, UEFA club competitions violated strategy-proofness mainly due to the unfair allocation of vacant slots. Vacant slots occur if teams can qualify for a tournament in more than one way, thus, the organiser needs to specify what happens if at least two conditions of qualification are satisfied. A straightforward example is the UEFA Europa League, where teams can enter via both their domestic championship and their domestic cup. According to Dagaev and Sonin, (2018), these tournament systems consisting of multiple round-robin and knockout tournaments are incentive incompatible if (1) there are at least two round-robin tournaments with at least one prize; (2) there is one round-robin tournament and at least one knockout tournament with at least one prize, and the vacant slots are not always filled through the round-robin tournament.
This theoretical finding might seem not especially interesting for sports policy. However, the misaligned allocation rule has been responsible for the incentive incompatibility of
  • the UEFA Europa League entry until the 2014/15 season (Dagaev and Sonin,, 2018);
  • the UEFA Champions League entry between the 2015/16 and 2017/18 seasons (Csató,, 2019);
  • the UEFA Champions League seeding from the 2015/16 season (Csató, 2020b, ).
Fortunately, Section 4 will reveal that the third problem has been solved by the new seeding rules effective from the 2024/25 season. On the other hand, the reform has also created potentially misaligned incentives in the UEFA Champions League qualification as will be seen in Section 3.
The introduction of the UEFA Nations League has also allowed for games where a team has an incentive to lose (Csató,, 2022; Haugen and Krumer,, 2021). Even though Scelles et al., (2024) identify no such games between 2018 and 2021, the authors call for more research around optimal competition formats that provide appropriate incentives for all teams in all games. Our paper is one of these follow-up studies.
3 A real-world example
Currently, the four highest-ranked associations—including Germany—are represented by four clubs in the UEFA Champions League. In addition, both the UEFA Champions League and UEFA Europa League titleholders automatically qualify for the league phase. The possible vacancy is filled according to UEFA, 2024b (, Article 3.04):
  1. a.
    A vacancy created by the UEFA Champions League titleholder is filled by the domestic champion with the highest individual club coefficient of all the clubs that qualify for the champions path.
  2. b.
    A vacancy created by the UEFA Europa League titleholder is filled by the club with the best individual club coefficient of all the clubs that qualify for the champions path and the league path, provided that this club is the highest ranked domestically of those from its association that have not already qualified for the league phase of the competition directly. If this condition of being the highest ranked domestic club is not met by the club with the highest individual club coefficient, then the position moves to the club with the next highest individual club coefficient of all clubs in the champions path and league path.
UEFA has introduced the so-called European performance spot from the 2024/25 season (UEFA, 2024b, , Article 3.07):
Finally, the two associations whose affiliated clubs achieved the best collective performance in the previous season of the UEFA men’s club competitions (i.e. best season club association coefficient), in accordance with the season coefficient principles (see Annex D.3), are each entitled to an additional place in the league phase, known as a European performance spot (EPS). The EPS is allocated to the club that finishes the relevant association’s domestic championship in the highest position of all those that do not qualify for the league phase of the UEFA Champions League via the domestic championship (after any vacancies have been filled in accordance with Articles 3.04 to 3.06). The two associations’ overall quota of clubs qualifying to UEFA club competitions are each increased by one club.
Table 1: (Hypothetical) standing of the 2023/24
German Bundesliga before the last matchday
Pos Team W D L GF GA GD Pts
1 Bayer Leverkusen 27 6 0 87 23 +++64 87
2 Bayern München 23 3 7 92 41 +++51 72
3 VfB Stuttgart 22 4 7 74 39 +++35 70
4 RB Leipzig 19 7 7 75 37 +++38 64
5 Borussia Dortmund 18 9 6 66 43 +++23 63
6 Eintracht Frankfurt 11 13 9 49 48 +++1 46
7 SC Freiburg 11 9 13 44 56 --12 42
8 1899 Hoffenheim 11 7 15 62 66 --4 40
  • The result of the match Borussia Dortmund vs. 1899 Hoffenheim has been changed from 2-3 to 4-3.
  • Pos = Position; W = Won; D = Drawn; L = Lost; GF = Goals for; GA = Goals against; GD = Goal difference; Pts = Points. All teams have played 33 matches.
Table 1 shows the standing of the 2023/24 German Bundesliga on the morning of 18 May 2024 (the last matchday) with a slight modification: the result of the match Borussia Dortmund vs. 1899 Hoffenheim, played on 15 February 2024, is changed from 2-3 to 4-3. At that moment, it was already known that Germany and Italy finished in the top two of the 2023/24 season association club coefficients, and earned the two European performance spots. The two clubs playing in the final of the UEFA Champions League on 1 June 2024 were known to be Borussia Dortmund and the Spanish club Real Madrid, too.
Consider the situation of Eintracht Frankfurt, which plays against RB Leipzig at home on the last (34th) matchday. It could be ranked neither higher nor lower than its current sixth position. Since the top four teams automatically qualify for the UEFA Champions League, the uncertainty with respect to this tournament resides in whether Borussia Dortmund or RB Leipzig obtains the fourth position, and whether Borussia Dortmund wins the 2023/24 UEFA Champions League or not. They determine the fate of Eintracht Frankfurt as follows:
  • If Borussia Dortmund is the fourth, then the first five teams qualify for the UEFA Champions League independently of which team wins the 2023/24 UEFA Champions League final. If Dortmund wins the final, a vacancy is created, filled by the Ukrainian champion Shakhtar Donetsk (UEFA, 2024b, , Article 3.04). If Dortmund loses the final, this team creates no vacancy.
  • If Borussia Dortmund is the fifth, then the first four teams automatically qualify for the UEFA Champions League. If Dortmund wins the 2023/24 UEFA Champions League final, it qualifies for the next season of this tournament as the titleholder. Thus, the European performance spot is obtained by the sixth-ranked Eintracht Frankfurt. If Dortmund loses the final, it qualifies for the next season of the UEFA Champions League due to the European performance spot earned by the Bundesliga.
In the first case, Eintracht Frankfurt could never play in the UEFA Champions League. However, in the second case, it can participate in this competition with a non-negligible probability (if Dortmund wins the 2023/24 UEFA Champions League). Consequently, the dominant strategy of Eintracht Frankfurt is to guarantee that Borussia Dortmund is ranked fifth in the 2023/24 German Bundesliga, which can be easily achieved by deliberately losing against RB Leipzig on 18 May 2024.
The above pathological situation has not been realised as Borussia Dortmund has been guaranteed to finish in the fifth place before the last round. Nonetheless, UEFA has clearly had good luck in avoiding a strange game in the 2023/24 German Bundesliga. Some football fans have also noticed the problem of misaligned incentives for Eintracht Frankfurt, see https://m.sports.ru/tribuna/blogs/gazzzzzpacho/3239387.html.
4 Discussion and policy implications
A recent literature survey on fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to accumulate more knowledge on the prevention of fraud. Our paper has examined the new format of the UEFA Champions League from this perspective. It should be acknowledged that the rules have improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for the guaranteed place of the titleholder in the first pot (UEFA, 2024b, , Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
On the other hand, strategy-proofness is violated in the new format according to Section 3. What is the underlying reason and how can incentive compatibility be restored?
UEFA, 2024b (, Article 3.04) has changed compared to the previous season of 2023/24. However, under both rules, no club could have an incentive to create a vacancy in order to obtain the released slot since either the highest-ranked not already qualified club from the association of the titleholder or a club from a different association benefits from the vacancy (see the cumbersome discussion of UEFA, 2024b (, Article 3.04b)). Consequently, Article 3.04 is not responsible for incentive incompatibility on its own.
On the other hand, the European performance spots are allocated only after the vacancies caused by the two titleholders are solved (UEFA, 2024b, , Article 3.07). Thus, a team playing in the same round-robin tournament (domestic championship) as one of the (potential) titleholders can be interested in not creating vacancies in the UEFA Champions League. This can be achieved by “helping” other teams to be ranked over the (potential) titleholder—a clear example of a tanking opportunity as we have seen above. Such misaligned incentives can arise in the associations entitled to at least one spot in the league phase of the UEFA Champions League if they also earn one of the European performance spots. These are currently the 10 highest-ranked associations: in the 2024/25 season, England, Spain, Germany, Italy, France, the Netherlands, Portugal, Belgium, Scotland, and Austria.
We offer the following remedies:
  • Rule change A: Modify the order of applying Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA, 2024b, ). That is, first the two European performance spots are allocated (UEFA, 2024b, , Article 3.07), and then the possible vacancies created by the titleholders are addressed (UEFA, 2024b, , Article 3.04).
  • Rule change B: Modify the way of filling vacancies in UEFA, 2024b (, Article 3.04). That is, any vacancy created by a titleholder is filled by the highest-ranked not already qualified club from its association.
Both reforms satisfy incentive compatibility. For Rule change A, the argument is provided above. For Rule change B, strategy-proofness is implied by Dagaev and Sonin, (2018, Proposition 6).
In the example of Section 3, Rule change A means that the top five German teams enter the UEFA Champions League due to the European performance spot, and Borussia Dortmund would create a vacancy if it wins the 2023/24 UEFA Champions League, to be filled by the Ukrainian champion Shakhtar Donetsk. On the other hand, Rule change B implies that the top five German teams enter the UEFA Champions League due to the European performance spot, and Borussia Dortmund would create a vacancy if it wins the 2023/24 UEFA Champions League, to be filled by the sixth-placed Eintracht Frankfurt.
Since Rule change A is closer to the original intention of UEFA, this would be our first recommendation. Rule change B favours the domestic championship of the titleholder, hence, it is expected to lead to an even higher concentration of the clubs playing in the UEFA Champions League by their associations because the highest-ranked countries are more likely to win both the UEFA Champions League and the UEFA Europa League.
If UEFA is not much worried about manipulation in a domestic league, and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. Following the previous literature, the current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, (2022), even if it would be difficult to implement.
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig is played in the last round. The set of top teams can be relatively accurately estimated at the beginning of the season, hence, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule may lead to a tanking opportunity at the end of the season with a (much) lower probability. However, even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before. Therefore, it is the responsibility of UEFA to act if it is really committed to its announced principles.
5 Conclusions
Our paper has revealed that the UEFA Champions League rules are incentive incompatible again from the 2024/25 season. The order in which the vacancies caused by the titleholders are filled and the two European performance spots are allocated has almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. UEFA is strongly encouraged to adopt one of our recommendations in order to safeguard the integrity of its flagship tournament.
The issue presented above provides an instructive lesson for sports governing bodies around the world: commitment to sporting merit requires not only media releases but appropriately designed competitions. Since rule changes such as the introduction of the European performance spot can always affect fairness, a thorough assessment is necessary before they are implemented. In order to achieve this fundamental aim, Kendall and Lenten, (2017) suggest two alternative calls to action:
  • The governing bodies of major sports should invite academic representatives onto their committees, who would be tasked with identifying possible loopholes in proposed rule changes in consultation with the scientific community; or
  • All proposed rule changes should be posted in a public forum, where interested academics could comment, perhaps after running simulations.
Hopefully, UEFA and all other sports governing bodies will adopt one of these options in the future. This would be especially important because, as it has been highlighted in Section 2, UEFA repeatedly commits the same mistake in allocating vacant slots—while the fair solution has been provided by Dagaev and Sonin, (2018), and the first version of this paper is available since March 2013.
Acknowledgements
We are grateful to Dries Goossens and David van Bulck, who have organised the third Fairness in Sports workshop in Ghent on 11 June 2024.
One anonymous colleague provided valuable comments and suggestions on an earlier draft.
We are indebted to the Wikipedia community for summarising important details of the sports competitions discussed in the paper.
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838, and the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.