### v6

# Misaligned incentives in the post-2024 UEFA Champions League: A managerial failure

Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification for the post-2024 UEFA Champions League is not always based on sporting merit as we show. In particular, a match with misaligned incentives has barely been avoided in the 2023/24 German Bundesliga. Our mathematical model reveals that fairness can be guaranteed by filling all vacant slots via the domestic leagues and retaining the association constraint in the league phase draw. UEFA repeatedly commits the same mistake in treating vacancies, which indicates a clear managerial failure. Therefore, governing bodies in sports are called to work more closely together with the scientific community.
Keywords: fairness; incentive compatibility; integrity; tournament design; UEFA Champions League
MSC class: 62F07, 90B90, 91B14
JEL classification number: C44, D71, Z20
Hoc est, quod ait Heraclitus: “In idem flumen bis descendimus et non descendimus.”111 Source: https://www.perseus.tufts.edu/hopper/text?doc=Perseus%3Atext%3A2007.01.0080%3Aletter%3D58
[…]
1 Introduction
Ensuring the fairness and integrity of sports is a crucial responsibility of tournament organisers (Kendall and Lenten,, 2017; Devriesere et al.,, 2024). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League in 2024 with the following words (UEFA, 2024a, ): “UEFA has clearly shown that we are fully committed to respecting the fundamental values of sport and to defending the key principle of open competitions, with qualification based on sporting merit, fully in line with the values and solidarity-based European sports model.
However, is the UEFA Champions League qualification based on sporting merit under all circumstances? Is there any issue in its rules that threatens the fairness of the tournament and can potentially jeopardise competition?
The current paper uncovers that, while UEFA may be committed to sporting merit in principles, their regulations do not exclude an unfair game. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots is assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig. A mathematical model is presented to understand the issue and compare different allocation rules of vacant slots. Two possible solutions are offered to decrease the probability of such a problem or even entirely avoid it. It is also shown that the association constraint in the league phase draw has a role in improving fairness.
The novel case of incentive incompatibility is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, the root of the problem resides in the allocation of vacant slots, which led to some instances of incentive incompatibility in the past. Making the same mistake repeatedly seems to indicate a clear managerial failure.
2 Related literature
Since tanking—the act of deliberately losing a game—is against the spirit of sports, incentive compatibility is an extensively discussed topic in the field of tournament design (Kendall and Lenten,, 2017; Csató,, 2021). Probably the most serious type of incentive incompatibility is when a team can be strictly better off by losing. According to our knowledge, such a possibility has been identified first by Dagaev and Sonin, (2018) in the UEFA Europa League. Several qualification tournaments for the FIFA World Cups and the UEFA European Championships have been verified to suffer from the same weakness (Csató, 2020a, ), even allowing for a match where the dominant strategy of both opponents is to play a draw (Csató, 2020c, ).
In the past, UEFA club competitions violated strategy-proofness mainly due to the unfair allocation of vacant slots. Vacant slots occur if teams can qualify for a tournament in more than one way, thus, the organiser needs to specify what happens if two conditions of qualification are simultaneously satisfied. A straightforward example is the UEFA Europa League, where teams can enter via both their domestic championship and domestic cup. According to Dagaev and Sonin, (2018), these tournament systems consisting of multiple round-robin and knockout tournaments are incentive incompatible if (1) there are at least two round-robin tournaments with at least one prize; (2) there is one round-robin tournament and at least one knockout tournament with at least one prize, and the vacant slots are not always filled through the round-robin tournament.
This theoretical result might not seem especially interesting for sports policy. However, the misaligned allocation rule has been responsible for the incentive incompatibility of
  • the UEFA Europa League entry until the 2014/15 season (Dagaev and Sonin,, 2018);
  • the UEFA Champions League entry between the 2015/16 and 2017/18 seasons (Csató,, 2019);
  • the UEFA Champions League seeding from the 2015/16 season (Csató, 2020b, ).
Fortunately, the third problem has been solved by the new seeding rules effective from the 2024/25 season, see Section 5. On the other hand, the reform has also created potentially misaligned incentives in the UEFA Champions League qualification as will be shown in Section 3.
The introduction of the UEFA Nations League has also allowed for matches where a team has an incentive to lose (Csató,, 2022; Haugen and Krumer,, 2021). Even though Scelles et al., (2024) identify no such games between 2018 and 2021, the authors call for more research around optimal competition formats that provide appropriate incentives for all teams in all games. Our paper can be considered as one of these follow-up studies.
3 A real-world example
The four highest-ranked associations—including Germany—are currently represented by four clubs in the UEFA Champions League. In addition, both the UEFA Champions League and UEFA Europa League titleholders automatically qualify for the league phase. The possible vacancy is filled according to UEFA, 2024b (, Article 3.04):
  1. a.
    A vacancy created by the UEFA Champions League titleholder is filled by the domestic champion with the highest individual club coefficient of all the clubs that qualify for the champions path.
  2. b.
    A vacancy created by the UEFA Europa League titleholder is filled by the club with the best individual club coefficient of all the clubs that qualify for the champions path and the league path, provided that this club is the highest ranked domestically of those from its association that have not already qualified for the league phase of the competition directly. If this condition of being the highest ranked domestic club is not met by the club with the highest individual club coefficient, then the position moves to the club with the next highest individual club coefficient of all clubs in the champions path and league path.
UEFA has introduced the so-called European performance spot from the 2024/25 season (UEFA, 2024b, , Article 3.07):
Finally, the two associations whose affiliated clubs achieved the best collective performance in the previous season of the UEFA men’s club competitions (i.e. best season club association coefficient), in accordance with the season coefficient principles (see Annex D.3), are each entitled to an additional place in the league phase, known as a European performance spot (EPS). The EPS is allocated to the club that finishes the relevant association’s domestic championship in the highest position of all those that do not qualify for the league phase of the UEFA Champions League via the domestic championship (after any vacancies have been filled in accordance with Articles 3.04 to 3.06). The two associations’ overall quota of clubs qualifying to UEFA club competitions are each increased by one club.
Table 1: (Hypothetical) standing of the 2023/24
German Bundesliga before the last matchday
Pos Team W D L GF GA GD Pts
1 Bayer Leverkusen 27 6 0 87 23 $+$64 87
2 Bayern München 23 3 7 92 41 $+$51 72
3 VfB Stuttgart 22 4 7 74 39 $+$35 70
4 RB Leipzig 19 7 7 75 37 $+$38 64
5 Borussia Dortmund 18 9 6 66 43 $+$23 63
6 Eintracht Frankfurt 11 13 9 49 48 $+$1 46
7 SC Freiburg 11 9 13 44 56 --12 42
8 1899 Hoffenheim 11 7 15 62 66 --4 40
  • The result of the match Borussia Dortmund vs. 1899 Hoffenheim has been changed from 2-3 to 4-3.
  • Pos = Position; W = Won; D = Drawn; L = Lost; GF = Goals for; GA = Goals against; GD = Goal difference; Pts = Points. All teams have played 33 matches.
Table 1 shows the standing of the 2023/24 German Bundesliga on the morning of 18 May 2024 (the last matchday) with a slight modification: the result of the match Borussia Dortmund vs. 1899 Hoffenheim, played on 15 February 2024, is changed from 2-3 to 4-3. At that moment, it was already known that Germany and Italy finished in the top two of the 2023/24 season association club coefficients, and earned the two European performance spots. The two clubs playing in the final of the UEFA Champions League on 1 June 2024 were known to be Borussia Dortmund and the Spanish club Real Madrid, too.
Consider the situation of Eintracht Frankfurt, which plays against RB Leipzig at home on the last (34th) matchday. It could be ranked neither higher nor lower than its current sixth position. Since the top four teams automatically qualify for the UEFA Champions League, the uncertainty with respect to this tournament resides in whether Borussia Dortmund or RB Leipzig obtains the fourth position, and whether Borussia Dortmund wins the 2023/24 UEFA Champions League or not. They determine the fate of Eintracht Frankfurt as follows:
  • If Borussia Dortmund is the fourth, then the first five teams qualify for the UEFA Champions League independently of which team wins the 2023/24 UEFA Champions League final. If Dortmund wins the final, a vacancy is created, filled by the Ukrainian champion Shakhtar Donetsk (UEFA, 2024b, , Article 3.04). If Dortmund loses the final, this team creates no vacancy.
  • If Borussia Dortmund is the fifth, then the first four teams automatically qualify for the UEFA Champions League. If Dortmund wins the 2023/24 UEFA Champions League final, it qualifies for the next season of this tournament as the titleholder. Thus, the European performance spot is obtained by the sixth-ranked Eintracht Frankfurt. If Dortmund loses the final, it qualifies for the next season of the UEFA Champions League due to the European performance spot earned by the Bundesliga.
In the first case, Eintracht Frankfurt could never play in the UEFA Champions League. However, in the second case, it can participate in this competition with a non-negligible probability (if Dortmund wins the 2023/24 UEFA Champions League). Consequently, the dominant strategy of Eintracht Frankfurt is to guarantee that Borussia Dortmund is ranked fifth in the 2023/24 German Bundesliga, which can be easily achieved by deliberately losing against RB Leipzig on 18 May 2024.
The above pathological situation has not been realised as Borussia Dortmund has been guaranteed to finish in the fifth place before the last round. Nonetheless, UEFA has clearly had good luck in avoiding a strange game in the 2023/24 German Bundesliga. Some football fans have also noticed the problem of misaligned incentives for Eintracht Frankfurt, see https://m.sports.ru/tribuna/blogs/gazzzzzpacho/3239387.html.
4 A mathematical model
A tournament $(X,Y)$ is a tuple of a set of results $X$ and a ranking method $Y$ that assigns a linear order of the teams to any set of results $X$. The number of teams in tournament $(X,Y)$ is denoted by $\lvert X\rvert$. For each team $1\leq i\leq\lvert X\rvert$, $Y_{i}(X)$ is its rank under a given set of results $X$.
Let $\mathbf{W}=\left[(S,P),(T,Q),(U,R)\right]$ be a tournament system consisting of three tournaments $(S,P)$, $(T,Q)$, $(U,R)$. The winner of tournament $(S,P)$ is denoted by $p$, that is, $1\leq p\leq\lvert S\rvert$ is the (unique) team for which $P_{p}(S)=1$ holds. This team $p$ also plays in tournament $(T,Q)$ but not in tournament $(U,R)$. The winner of tournament $(U,R)$ is denoted by $r$: $1\leq r\leq\lvert U\rvert$ is the (unique) team for which $R_{r}(U)=1$ holds.
This model can represent the qualification of the UEFA Champions League as follows. $(S,P)$ is the Champions League. $(T,Q)$ is the domestic league of the Champions League winner, denoted by $p$. $(U,R)$ is the “competition” played by all teams qualified for the champions path in the qualification of the Champions League such that the ranking method $R$ is based on the club coefficients of these teams.
The allocation rule $\mathcal{A}(\mathbf{W})$ chooses the winner of tournament $(S,P)$ and the top $q$ teams of tournament $(T,Q)$. However, a vacancy may be created if team $p$ is ranked among the top $q$ teams in tournament $(T,Q)$.
We consider three allocation rules that differ in the assignment of the vacant slot caused by team $p$. As discussed in Section 3, the current rule of the UEFA Champions League is:
$\mathcal{A}(\mathbf{W})=\begin{cases}i:i=p\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\geq q+1\\ i:i=r\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\leq q-1\\ i:Q_{i}(T)\leq q+1&\text{if }Q_{p}(T)=q.\end{cases}$
Note that the winner $r$ of tournament $(U,R)$ receives the vacant slot only if the titleholder team $p$ is ranked among the top $q-1$ teams in tournament $(T,Q)$ because the last, $q$th slot of tournament $(T,Q)$ is the European performance spot.
The allocation rule is simplified if the order of Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA, 2024b, ) is reversed. Then the two European performance spots are allocated first (UEFA, 2024b, , Article 3.07), followed by addressing the possible vacancies created by the titleholders (UEFA, 2024b, , Article 3.04), which leads to:
$\mathcal{A}^{\prime}(\mathbf{W})=\begin{cases}i:i=p\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\geq q+1\\ i:i=r\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\leq q.\end{cases}$
Finally, any vacancy created by a titleholder can be filled by the highest-ranked not already qualified club from its association:
$\mathcal{A}^{\prime\prime}(\mathbf{W})=\begin{cases}i:i=p\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\geq q+1\\ i:Q_{i}(T)\leq q+1&\text{if }Q_{p}(T)\leq q.\end{cases}$
Now the incentive (in)compatibility of these allocation rules can be explored.
Definition 1.
Tournament $(S,P)$ is called manipulable by team $1\leq i\leq\lvert S\rvert$ if there exists a tournament $\left(S^{\prime},P\right)$ such that $S^{\prime}=S$ except for a worse performance by team $i$, and the following two conditions hold:
  • $Q_{i}\left(S\right)=Q_{i}\left(S^{\prime}\right)$;
  • $Q_{j}\left(S\right)\geq 2$ but $Q_{j}\left(S^{\prime}\right)=1$ for a team $1\leq j\leq\lvert S\rvert$.
That is, a tanking strategy by team $i$ guarantees the same final ranking for team $i$ and makes team $j$ the tournament winner.
Definition 2.
Tournament $(T,Q)$ is called manipulable by team $1\leq i\leq\lvert T\rvert$ if there exists a tournament $\left(T^{\prime},Q\right)$ such that $T^{\prime}=T$ except for a worse performance by team $i$, and the following two conditions hold:
  • $Q_{i}\left(T\right)=Q_{i}\left(T^{\prime}\right)=q+1$;
  • $Q_{j}\left(T\right)=Q_{k}\left(T^{\prime}\right)=q-1$, as well as $Q_{j}\left(T^{\prime}\right)=Q_{k}\left(T\right)=q$ for some teams $1\leq j,k\leq\lvert T\rvert$.
That is, a tanking strategy by team $i$ guarantees that it remains ranked $(q+1)$th, while reverses teams $j$ and $k$ in the two positions directly above.
In tournament $(U,R)$, no tanking strategy exists because the UEFA club coefficients cannot be increased by a worse performance on the field.
Proposition 1.
Tournament system $\mathbf{W}=\left[(S,P),(T,Q),(U,R)\right]$ with the allocation rule $\mathcal{A}(\mathbf{W})$ is incentive incompatible if and only if one of the following conditions hold:
  • tournament $(S,P)$ is manipulable by the team $1\leq r\leq\lvert U\rvert$ (for which $R_{r}(U)=1$) such that the team $j$ in Definition 1 plays in both tournaments $(S,P)$ and $(T,Q)$ with $Q_{j}(T)\leq q-1$; or
  • tournament $(S,P)$ is manipulable by the team $1\leq i\leq\lvert T\rvert$ for which $Q_{i}(T)=q+1$ such that the team $j$ in Definition 1 plays in both tournaments $(S,P)$ and $(T,Q)$ with $Q_{j}(T)=q$; or
  • tournament $(T,Q)$ is manipulable by the team $1\leq i\leq\lvert T\rvert$ for which $Q_{i}(T)=q+1$ such that the team $j$ in Definition 2 plays in both tournaments $(S,P)$ and $(T,Q)$ with $P_{j}(S)=1$, that is, $j=p$.
Proof.
If the first condition is satisfied, team $r$ can create a vacancy by making team $j$ the winner of tournament $(S,P)$, which is filled by team $r$ since $Q_{j}(T)\leq q-1$.
If the second condition is satisfied, team $i$ can create a vacancy by making team $j$ the winner of tournament $(S,P)$, which is filled by team $i$ since $Q_{j}(T)=q$.
If the third condition is satisfied, team $i$ can create a vacancy by achieving that the winner of tournament $(S,P)$ is ranked $q$th in tournament $(T,Q)$, which is filled by team $i$.
The definition of the allocation rule $\mathcal{A}$ implies that no other team can receive the possible vacant slot, thus, incentive incompatibility can arise only in the cases above. ∎
The example presented in Section 3 exploits the third condition.
Proposition 2.
Tournament system $\mathbf{W}=\left[(S,P),(T,Q),(U,R)\right]$ with the allocation rule $\mathcal{A}^{\prime}(\mathbf{W})$ is incentive incompatible if and only if tournament $(S,P)$ is manipulable by the team $1\leq r\leq\lvert U\rvert$ (for which $R_{r}(U)=1$) such that the team $j$ in Definition 1 plays in both tournaments $(S,P)$ and $(T,Q)$ with $Q_{j}(T)\leq q$.
Proof.
It follows from the proof of Proposition 1. The second and third conditions in Proposition 1 are eliminated because all cases $Q_{p}(T)\leq q$ are treated uniformly by the allocation rule $\mathcal{A}^{\prime}$. ∎
Proposition 3.
Tournament system $\mathbf{W}=\left[(S,P),(T,Q),(U,R)\right]$ with the allocation rule $\mathcal{A}^{\prime\prime}(\mathbf{W})$ is incentive incompatible if and only if tournament $(S,P)$ is manipulable by the team $1\leq i\leq\lvert T\rvert$ for which $Q_{i}(T)=q+1$ such that the team $j$ in Definition 1 plays in both tournaments $(S,P)$ and $(T,Q)$ with $Q_{j}(T)=q$.
Proof.
It follows from the proof of Proposition 1. The first and third conditions in Proposition 1 are eliminated because all cases $Q_{p}(T)\leq q$ are treated uniformly by the allocation rule $\mathcal{A}^{\prime\prime}$. ∎
The results above cannot be derived from the model of Dagaev and Sonin, (2018) since the latter contains multiple round-robin and knockout tournaments. However, if $(S,P)$ is a knockout tournament, then it cannot be manipulated by any team $1\leq i\leq\lvert S\rvert$, and both allocation rules $\mathcal{A}^{\prime}$ and $\mathcal{A}^{\prime\prime}$ guarantee incentive compatibility.
5 Discussion
A recent survey on fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to accumulate more knowledge on the prevention of fraud. Our paper examines the new format of the UEFA Champions League from this perspective. The rules have clearly improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for the guaranteed place of the titleholder in the first pot (UEFA, 2024b, , Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
On the other hand, strategy-proofness is violated in the new format according to Section 3. The underlying reason is highlighted by the analysis of the mathematical model in Section 4, in particular, by Proposition 1: the European performance spots are allocated only after the vacancies caused by the two titleholders are solved (UEFA, 2024b, , Article 3.07). Thus, a team playing in the same round-robin tournament (domestic league) as one of the (potential) titleholders is interested in not creating vacancies in the UEFA Champions League. This can be achieved by “helping” other teams to be ranked over the (potential) titleholder. Such misaligned incentives may arise in the associations entitled to at least one spot in the league phase of the UEFA Champions League if they also earn one of the European performance spots. These are currently the 10 highest-ranked associations: in the 2024/25 season, England, Spain, Germany, Italy, France, the Netherlands, Portugal, Belgium, Scotland, and Austria.
Based on Section 4, incentive incompatibility is mitigated by changing the allocation rule according to the definition of either $\mathcal{A}^{\prime}$ or $\mathcal{A}^{\prime\prime}$:
  • Rule change A: Modify the order of applying Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA, 2024b, ). That is, first the two European performance spots are allocated (UEFA, 2024b, , Article 3.07), and then the possible vacancies created by the titleholders are addressed (UEFA, 2024b, , Article 3.04).
  • Rule change B: Modify the way of filling vacancies in UEFA, 2024b (, Article 3.04). That is, any vacancy created by a titleholder is filled by the highest-ranked not already qualified club from its association.
If rule change A is chosen, Proposition 2 remains valid. Hence, a tanking strategy can exist in the league phase of the UEFA Champions League for a team that qualifies for the champions path in the qualification of the next Champions League. For example, any team that is already eliminated in the league phase but can have the highest UEFA club coefficient in the champions path is interested in creating a vacant slot. This might be achieved by losing against a team that is expected to obtain a direct qualifying slot in its domestic league in order to help it win the Champions League. Consider the Swiss champion Young Boys in the 2024/25 UEFA Champions League. The club was eliminated in the league phase after six rounds and played against Celtic (Scotland) and Crvena Zvezda (Serbia) in the last two rounds. Since the Scottish champion directly qualifies for the league phase in contrast to the Serbian champion, Young Boys could have maximised its chances by losing against Celtic and exerting full effort against Crvena Zvezda.
On the other hand, if rule change B is chosen, Proposition 3 should be considered. Hence, a tanking strategy can exist in the league phase of the UEFA Champions League for a team if the Champions League is won by a team from the same national association due to the tanking. However, this is impossible because two teams from the same country are not allowed to play against each other in the league phase (UEFA, 2024b, , Article 16.02), and a deliberate loss in the knockout stage cannot be beneficial.
Consequently, even though rule change A seems to be closer to the original intention of UEFA, it might still lead to incentive incompatibility, albeit with a lower probability. At least, the situation presented in Section 3 will certainly be avoided. In contrast, rule change B entirely avoids misaligned incentives and highlights the role of the association constraint in the league phase draw, which is desirable not only to maximise the number of international matches played but also to ensure strategy-proofness in the league phase. This is a good illustration of the idea raised by Csató, (2022) that draw restrictions can be used to prohibit potentially unfair matches. Nonetheless, rule change B favours the domestic league of the titleholder, hence, it leads to a somewhat higher concentration of the clubs playing in the UEFA Champions League by their associations because the highest-ranked countries are more likely to win both the UEFA Champions League and the UEFA Europa League.
To summarise, the main message of Dagaev and Sonin, (2018) also applies to the model of Section 4: the vacant slots should be filled via the round-robin tournament as done by the allocation rule $\mathcal{A}^{\prime\prime}$. Therefore, UEFA could have guaranteed incentive compatibility by following the recommendation of Dagaev and Sonin, (2018) despite the different settings.
In the example of Section 3, Rule change A means that the top five German teams enter the UEFA Champions League due to the European performance spot, and Borussia Dortmund would create a vacancy if it wins the 2023/24 UEFA Champions League, to be filled by the Ukrainian champion Shakhtar Donetsk. On the other hand, Rule change B implies that the top five German teams enter the UEFA Champions League due to the European performance spot, and Borussia Dortmund would create a vacancy if it wins the 2023/24 UEFA Champions League, to be filled by the sixth-placed Eintracht Frankfurt.
If UEFA is not worried much about manipulation in a domestic league and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. The current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, (2022), even if it would be difficult to implement.
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig was played in the last round. The set of top teams may be relatively accurately estimated at the beginning of the season, thus, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule might lead to a tanking opportunity at the end of the season with a (much) lower probability. But even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before. Therefore, it is the responsibility of UEFA to act if it is really committed to its announced principles.
6 Conclusions
Our paper has revealed that the UEFA Champions League rules are incentive incompatible again from the 2024/25 season. In particular, the allocation rules of the vacancies caused by the titleholders and the two European performance spots have almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. UEFA is strongly encouraged to adopt our recommendation in order to safeguard the integrity of its flagship tournament.
Table 2: Timeline of incentive incompatibility in UEFA club competitions
caused by the allocation of vacant slots
Date/Period Problem
2009/10–2014/15 Incentive incompatibility of the UEFA Europa League entry (Dagaev and Sonin,, 2018)
May 2012 The match Feyenoord vs. Heerenveen, where Heerenveen was better off by losing than by playing a draw (Csató,, 2021, Chapter 2.1)
March 2013 The first version of Dagaev and Sonin, (2018) becomes available that verifies the incentive incompatibility of the UEFA Europa League entry and offers a general solution to the problem
2015/16–2017/18 Incentive incompatibility of the UEFA Champions League entry (Csató,, 2019)
2015/16–2023/24 Incentive incompatibility of the UEFA Champions League seeding system (Csató, 2020b, )
2024/25– Incentive incompatibility of the UEFA Champions League entry (this paper)
The problem identified is especially worrying because it is rooted in the allocation of vacant slots, a topic that was repeatedly responsible for incentive incompatibility in UEFA club competitions as summarised in Table 2. This issue offers an instructive lesson for decision-makers around the world: commitment to sporting merit requires not only media releases but appropriately designed tournaments. Since rule changes such as the introduction of the European performance spot can always affect fairness, a thorough assessment is necessary before they are implemented.
In order to achieve that fundamental aim, Kendall and Lenten, (2017) suggest two alternative calls to action:
  • The governing bodies of major sports should invite academic representatives onto their committees, who would be tasked with identifying possible loopholes in proposed rule changes in consultation with the scientific community; or
  • All proposed rule changes need to be posted in a public forum, where interested academics could comment, perhaps after running simulations.
Hopefully, most sports governing bodies, including UEFA, will adopt one of these options in the near future.
Acknowledgements
We are grateful to Dries Goossens and David van Bulck, who have organised the third Fairness in Sports workshop in Ghent on 11 June 2024.
Phil Scarf and one anonymous colleague provided valuable comments and suggestions on an earlier draft.
We are indebted to the Wikipedia community for summarising important details of the sports competitions discussed in the paper.
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838, and the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.